OpenStudy (anonymous):

Will give a medal Two cans of vegetables were measured. Can A is 10 centimeters in diameter and has a height of 10.5 cm. Can B has a height of 11 cm and a diameter of 7.5 cm. Which can has the greater volume? A) Can A B)Can B C)The two cans have equal volumes. D)More information is needed to solve the problem.

OpenStudy (anonymous):

Can A

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

can you explain this and hi

OpenStudy (misty1212):

volume is \(\pi r^2h\) and \(r\) is half the diameter

OpenStudy (misty1212):

first on has volume \[\pi \times 5^2\times 10.5\]

OpenStudy (misty1212):

second one has volume \[\pi \times (3.75)^2\times 11\]

OpenStudy (misty1212):

since you are just comparing them you can ignore the \(\pi\) and compare \[5^2\times 10.5\] with \[(3.75)^2\times 11\]

OpenStudy (anonymous):

thank you